# Style-Adaptive Detection Transformer for Single-Source Domain Generalized Object Detection

### v1

@@ Abstract
Single-source Domain Generalization (SDG) in object detection aims to develop a detector using only data from source domain that can exhibit strong generalization capability when applied to other unseen target domains. Existing methods are built upon CNN-based detectors and primarily improve their robustness by employing carefully designed data augmentation strategies integrated with feature alignment techniques. However, the data augmentation methods have inherent drawbacks that are only effective when the augmented sample distribution approximates or covers the unseen scenarios, thus failing to enhance the detector’s generalization capability across all unseen scenarios. Furthermore, while the recent DEtection TRansformer (DETR) has demonstrated superior generalization capability in domain adaptation tasks due to its efficient global information extraction, its potential in SDG tasks still remains unexplored. To this end, we introduce a strong DETR-based detector named the Style-Adaptive DEtection TRansformer (SA-DETR) for SDG in object detection. Specifically, we present a domain style adapter to project the style representation of the unseen target domain into the training domain, enabling dynamic style adaptation across various unseen scenarios. Then, the object-aware contrastive learning module is proposed to force the detector to extract domain-invariant features of objects through contrastive learning. By employing carefully designed object-aware gating masks to constrain the scope of feature aggregation in both spatial position and semantic category, this module effectively achieves cross-domain contrast of instance-level features, thereby further enhancing the detector’s generalization capability. Extensive experimental results demonstrate the superior performance and generalization capability of our proposed SA-DETR across five different weather scenarios. Code is released at https://github.com/h751410234/SA-DETR.
@@ I Introduction
Recently, many methods[4, 5, 6] for SDG in object detection have been proposed, employing carefully designed generalization strategies to enhance the robustness of detectors, but they overlook two significant issues. Firstly, to overcome the limited sample space of single-source domain data, existing methods typically employ innovative data augmentation strategies (at the image or feature level) to obtain more diverse samples for training, thereby enhancing the generalization capability of detectors for unseen scenarios. However, these methods have inherent drawbacks that are only effective when the augmented sample distribution approximates or covers the unseen scenarios. Due to the unpredictable deployment environments, which are likely to have significant differences from the augmented samples, the improvement in detector performance is minimal in such cases.
@@
Secondly, researchers have primarily explored and demonstrated effectiveness within the Faster R-CNN[7], while transformer-based detectors have been less studied. Recent works[8, 9, 10] increasingly shows that detectors based on the DEtection TRansformer (DETR)[11] exhibit superior performance in handling domain gaps compared to those based on Convolutional Neural Networks (CNNs). Unlike purely convolutional designs, DETR innovatively integrates a CNN backbone with transformer architecture (i.e., encoder and decoder), leveraging the attention mechanism to refine features for more detailed representations[12]. Additionally, the transformer architecture has been proven to be highly effective in capturing global structural information [13, 14], potentially offering better generalization capability for unseen scenarios. Therefore, it is highly meaningful to explore the effectiveness of DETR-based detectors in addressing the SDG task.
In this paper, we introduce a strong DETR-based detector named the Style-Adaptive DEtection TRansformer (SA-DETR) for SDG in object detection. To better adapt to unseen scenarios, we designed a domain style adapter that projects the style representation of the unseen target domain into the training domain. Unlike methods that expand data space, our adapter establishes a new avenue that utilizes known data to represent unseen ones, thereby enabling dynamic style adaptation across various unseen scenarios. Specifically, we use a memory module to store the statistical data (mean and variance) of feature maps extracted from the backbone during each training iteration. These data are naturally considered to be prototypical features of image styles, similar to those used in style transfer techniques[15, 16, 17]. As the stored prototypes accumulate, the memory module can develop dataset-level style representations, which we use as the style basis to rectify the style of the unseen images. During inference, we first calculate the Wasserstein distance [18] between the statistical data of the input image and the stored prototypes to estimate the style distribution discrepancy. Then, this distance is used as weighting factors, integrated within the AdaIN function[19] to facilitate an accurate projection.
@@
  • We present a novel domain style adapter which projects the style representation of the unseen target domain into the training domain, enabling dynamic style adaptation across various unseen scenarios. The adapter establishes a new avenue for addressing the challenge of single-source domain generalization in object detection.
  • To further enhance the generalization capability, an object-aware contrastive learning module is introduced. This module employs carefully designed object-aware gating masks to constrain the scope of feature aggregation in both spatial position and semantic category, thus forcing the detector to extract domain-invariant features of objects.
Extensive results demonstrate the superior performance and generalization capability of SA-DETR compared to state-of-the-art methods across five different weather conditions. Our results consistently achieved the best outcomes in all scenarios and showed a remarkable $12.2\%$ improvement in mAP on the ”Dusk-Rainy” scene.
@@ II Related Work
II-A The Preliminaries of Detection Transformer
Following the success of the Transformer architecture[20], Carion et al.[11] integrated this architecture into an object detection method named DEtection TRansformer (DETR). Unlike traditional convolutional designs, DETR uses encoders to further refine features extracted from the backbone network and employs multiple object queries to obtain detection results in the decoder. This approach significantly simplifies the detection pipline by eliminating the need for manually designed anchors[7] and Non-Maximum Suppression (NMS)[21]. Recently, many researchers[22, 23, 24] have investigated the potential of DETR-based detectors in addressing domain adaptation challenges, given that the transformer architecture excels at capturing global structural information. However, for domain generalization in object detection, existing methods still build on fully convolutional two-stage architectures like Faster R-CNN[7]. In this paper, we showcase the superior generalization capability of DETR-based detectors in unseen scenarios and introduce a style-adaptive detection transformer. To the best of our knowledge, this is the first study to explore the benefit of a DETR-based detector for the single-source domain generalization task.
II-B Domain Adaptation and Generalization
@@
II-C Single-source Domain Generalization for Object Detection
@@
Single-source Domain Generalization (SDG), a specific case of domain generalization task, aims to enhance network performance on unseen scenarios using only data from one source domain. At present, most research on SDG has concentrated on image classification[28, 29, 30] and segmentation[31, 32, 33], with limited work in the field of object detection, which warrants further exploration. Existing approaches primarily enhance the detector’s generalization capability by implementing data augmentation strategies and feature alignment techniques.
@@
For data augmentation strategies, the methods are similar to those used in image classification. IBN-Net[34] and ISW[35] utilize instance normalization and whitening on extracted features to mitigate the impact of specific domain styles. Then, CLIP-Gap[5] employs the pre-trained vision-language model CLIP[36], using carefully designed text prompts to improve the semantic style features extracted from the backbone. Additionally, MAD[37] transfers input images to the frequency domain and randomizes the distribution of extremely high and low-frequency components to increase the diversity of source domain data. Building on this, UFR[6] incorporates SAM[38] to isolate targets within images for further augmentation. Unlike the aforementioned methods, our proposed domain style adapter projects the style representations of the unseen target domain into the training domain, enabling dynamic style adaptation across various unseen scenarios.
@@
Feature alignment techniques are usually applied in SDG for object detection to further alleviate the influence of the domain gap. These techniques align cross-domain features by utilizing data from both the source and augmented domains, thereby facilitating the extraction of domain-invariant features in the detector. SDGOD[4] introduces a cyclic-disentangled self-distillation framework that achieves domain-invariant feature extraction without annotated labels. MAD[37] proposes a multi-view adversarial discriminator to achieve more precise alignment of cross-domain features. To complement the image-level feature alignment methods discussed above, UFR[6] has developed a causal prototype learning module that refines features extracted from Regions Of Interest (ROI) to further mitigate the influence of confounding factors in object attributes. Since the proposed adapter effectively addresses the image-level domain gap, we introduce an object-aware contrastive learning module to further encourage the DETR-based detector to extract domain-invariant instance-level features. This module employs carefully designed object-aware gating masks to constrain contrastive learning on instance-level features, thereby achieving accurate cross-domain feature alignment for objects.
@@ III Method
Figure 1: Details of (a) the overview of the proposed SA-DETR framework, which adopts DINO as the base detector and incorporates a Domain Style Adapter (DSA) and an Object-Aware Contrastive Learning (OCL) module, and (b) the workflow of the DSA. Meanwhile, the memory module is only used during training to store statistical data of input features and to form a style base, as indicated by the blue arrows. (c) illustrating the projection process during inference, where red and blue respectively represent the source and augmented domains, and green represents the unseen target domain. The adapter rectifies the channel statistical characteristics of the backbone’s extracted feature maps using the stored style base, effectively projecting the style representations of the unseen target domain into the training domain.
@@ III-A Framework Overview
DETR revisit. In our approach, we utilize DINO[39] as the base detector, an advanced DETR-based end-to-end object detector. Specifically, DINO incorporates a CNN backbone for feature extraction and employs an encoder-decoder transformer structure along with a detection head for the final detection predictions. When processing an image, the backbone initially extracts multi-scale feature maps similar to those used in CNN-based detectors. These feature maps are then reshaped into feature tokens that feed into the encoder-decoder module. The encoder further refines the input features by using self-attention. The decoder uses inserted object queries to probe local regions and aggregate instance features from the refined features through cross-attention, generating output embeddings. Finally, these output embeddings are used to compute prediction results through the detection head.
@@
Data augmentation. The objective of single-source domain generalization is to train the network with data from a single domain and to generalize its performance to a variety of unseen scenarios. In this task, data augmentation methods are essential to break the limitations of sample diversity. Inspired by previous work[37], we employ spurious correlation generator to augment the source dataset at the image level, ultimately yielding an augmented dataset of equivalent size. Therefore, the input to our method includes both the source and augmented data. Building upon this foundation, we propose the domain style adapter to leverage known styles for representing unseen ones, thereby overcoming the limitations of augmentation methods in improving performance.
@@
Framework. As shown in Fig. 1 (a), Our SA-DETR adopts DINO as the base detector and integrates our designed methods, including a Domain Style Adapter (DSA) and an Object-aware Contrastive Learning (OCL) module. Specifically, during training, the network inputs consist of pairs of images, each pair containing an equal number of images from both the source and augmented domains. For a batch of image pairs, the detector first uses a ResNet backbone[40] to extract multi-scale feature maps, and feeds the maps from each layer into our proposed adapter. The adapter stores the statistical data of the input features using a memory module and forms them into two domain-style bases. Then, the extracted feature maps are further refined in the encoder. Meanwhile, we insert an OCL module into the encoder, which utilizes contrastive learning to force the extraction of domain-invariant features of objects, thereby enhancing the detector’s generalization capability. Ultimately, the refined features are input into the subsequent decoder and detection head to obtain detection results, in line with DINO. During inference, the adapter rectifies the channel statistical characteristics of the backbone’s extracted feature maps using the stored style base, effectively projecting the style representations of the unseen domain into the training domain. Notably, our OCL module is employed only during training as a plugin. More details on the DSA and the OCL module will be presented in Subsections III-B and III-C, respectively.
III-B Domain Style Adapter
Previous works have adopted data augmentation to increase dataset diversity and then trained on the mixed data to generalize the network. However, these methods are effective only when the augmented sample distributions closely match the unseen scenarios, as detailed in the experimental section IV-C. Inspired by style transfer techniques[16, 31, 41], we propose a domain style adapter that projects the style representation of the target domain into the training domain for dynamic style adaptation across various unseen scenarios.

### v2

@@ Abstract
Single-source Domain Generalization (SDG) in object detection aims to develop a detector using only data from source domain that can exhibit strong generalization capability when applied to other unseen target domains. Existing methods are built upon CNN-based detectors and primarily improve their robustness by employing carefully designed data augmentation strategies integrated with feature alignment techniques. However, the data augmentation methods have inherent drawbacks that are only effective when the augmented sample distribution approximates or covers the unseen scenarios, thus failing to enhance the detector’s generalization capability across all unseen scenarios. Furthermore, while the recent DEtection TRansformer (DETR) has demonstrated superior generalization capability in domain adaptation tasks due to its efficient global information extraction, its potential for SDG tasks remains unexplored. To this end, we introduce a strong DETR-based detector named the Style-Adaptive DEtection TRansformer (SA-DETR) for SDG in object detection. Specifically, we present an online domain style adapter to project the style representation of the unseen target domain into the training domain. This adapter maintains a dynamic memory bank that self-organizes onto diverse style bases and continuously absorbs information from unseen domains within a test-time adaptation framework, ultimately delivering precise style adaptation. Then, the object-aware contrastive learning module is proposed to force the detector to extract domain-invariant features of objects through contrastive learning. By employing carefully designed object-aware gating masks to constrain the scope of feature aggregation in both spatial position and semantic category, this module effectively achieves cross-domain contrast of instance-level features, thereby further enhancing the detector’s generalization capability. Extensive experimental results demonstrate the superior performance and generalization capability of our proposed SA-DETR across five different weather scenarios. Code is released at https://github.com/h751410234/SA-DETR.
@@ I Introduction
Recently, numerous single-source domain generalization (SDG) methods [4, 5, 6] for object detection have been proposed, employing carefully designed generalization strategies to enhance detector robustness, yet they still suffer from two major limitations. First, to overcome the limited sample space of single-domain data, existing methods typically employ innovative data augmentation strategies (at the image or feature level) to obtain more diverse samples for training, thereby enhancing the generalization capability of detectors for unseen scenarios. These methods have inherent drawbacks that are only effective when the augmented sample distribution approximates or covers the unseen scenarios. Due to the unpredictable deployment environments, which are likely to have significant differences from the augmented samples, the improvement in detector performance is minimal in such cases. Second, another line of research tackles this issue through style modulation, suppressing domain‑specific styles via instance normalization [7] and feature whitening [8], or rectifying input images with simply constructed style bases [9, 10]. Such a static treatment overlooks the high diversity and dynamic nature of styles in the wild. Consequently, it remains insufficiently adaptive when confronted with large domain gaps.
@@
Moreover, researchers have primarily explored and demonstrated effectiveness within the Faster R-CNN[11], while transformer-based detectors have been less studied. Recent works[12, 13, 14] increasingly shows that detectors based on the DEtection TRansformer (DETR)[15] exhibit superior performance in handling domain gaps compared to those based on Convolutional Neural Networks (CNNs). Unlike purely convolutional designs, DETR innovatively integrates a CNN backbone with transformer architecture (i.e., encoder and decoder), leveraging the attention mechanism to refine features for more detailed representations[16]. Additionally, the transformer architecture has been proven to be highly effective in capturing global structural information [17, 18, 19], potentially offering better generalization capability for unseen scenarios. Therefore, it is highly meaningful to explore the effectiveness of DETR-based detectors in addressing the SDG task.
In this paper, we introduce a strong DETR-based detector named the Style-Adaptive DEtection TRansformer (SA-DETR) for SDG in object detection. To better adapt to unseen scenarios, we propose the Online Domain Style Adapter (ODS-Adapter), which addresses the domain gap through statistical matching rather than data augmentation. During training, the adapter employs a dynamic style memory bank to cache the channel-wise statistics (mean and variance) of backbone features and treats them as style bases, consistent with the classic style-transfer paradigm [20]. As iterations progress, similar bases are integrated through momentum-based updates, whereas redundant ones are discarded, allowing the memory bank to self-organize into multiple style bases that faithfully capture the diversity of the training data. During inference, the method first employs the Wasserstein distance[21] to quantify distribution discrepancies between the statistics of the testing image and the stored style bases. Next, we convert the distances into weights, which are injected into an AdaIN-like modulation to achieve fast and precise feature projection. Considering that the target domain may exhibit entirely novel styles, the adapter writes the newly observed statistics back to the memory bank during testing, thereby continuously absorbing information from unseen domains and preventing under-adaptation caused by large style gaps.
@@
  • We introduce an Online Domain Style Adapter (ODS-Adapter) that projects the style representation of unseen target domains onto the training domain. By maintaining a dynamic memory bank that self-organizes into diverse style bases and continuously absorbs statistics from unseen domains under a test time adaptation framework, the adapter enables rapid and accurate style adaptation to previously unseen scenarios.
  • To further enhance the generalization capability, an Object-aware Contrastive Learning (OCL) module is introduced. This module employs carefully designed object-aware gating masks to constrain the scope of feature aggregation in both spatial position and semantic category, thus forcing the detector to extract domain-invariant features of objects.
[…]
@@ II Related Work
II-A Domain Adaptation and Generalization
@@
II-B Single-source Domain Generalization for Object Detection
@@
Single-source Domain Generalization (SDG), a specific case of domain generalization task, aims to enhance network performance on unseen scenarios using only data from one source domain. At present, most research on SDG has concentrated on image classification[25, 26, 27, 28] and segmentation[29, 30, 31], with limited work in the field of object detection, which warrants further exploration. Existing approaches primarily enhance the detector’s generalization capability by implementing data augmentation strategies and feature alignment techniques.
@@
Data augmentation strategies improve model robustness by perturbing images or features to simulate variations that may occur in unseen scenarios. MAD [32] transfers input images into the frequency domain and randomizes the distribution of extremely high- and low-frequency components to increase the diversity of source domain data. Building on this, UFR [6] incorporates SAM [33] to isolate foreground objects from the background, allowing for targeted augmentation of each component. Liu et al. [34] further applies perturbations at both the image and feature levels to simulate a range of color and style biases. CLIP the Gap [5] leverages the pre-trained vision-language model CLIP [35], using carefully crafted textual prompts to simulate unknown scene styles and thereby enhance the semantic style features extracted by the backbone. PGST [36] improves the alignment between textual and visual features by leveraging GLIP [37] to project the visual style of source image regions into a hypothetical target domain style. In contrast to the above methods, our proposed adapter projects the style representation of the unseen domain onto the training domain, enabling dynamic style adaptation across diverse unseen scenarios rather than being constrained by the limited sample space of augmented training data.
@@
Feature alignment techniques are usually applied in SDG for object detection to further alleviate the influence of the domain gap. These techniques align cross-domain features by utilizing data from both the source and augmented domains, thereby facilitating the extraction of domain-invariant features in the detector. SDGOD [4] introduces a cyclic-disentangled self-distillation framework that achieves domain-invariant feature extraction without annotated labels. MAD [32] proposes a Multi-view Adversarial Discriminator to achieve more precise alignment of cross-domain features. To complement the image-level feature alignment methods discussed above, UFR [6] has developed a causal prototype learning module that refines features extracted from regions of interest (ROIs) to further mitigate the influence of confounding factors in object attributes. Since the proposed adapter effectively addresses the image-level domain gap, we introduce an object-aware contrastive learning module to further encourage the DETR-based detector to extract domain-invariant instance-level features. This module creates object-aware gating masks to constrain contrastive learning on instance-level features, thereby achieving accurate cross-domain feature alignment for objects.
II-C Style Modulation for Domain Adaptation
Unlike the aforementioned data augmentation approaches, modulating the statistical properties of features provides an effective alternative for style adaptation and has been widely adopted in domain adaptation and generalization tasks. IBN-Net [7] and ISW [8] first introduced instance normalization and whitening operations to mitigate the influence of domain-specific styles on feature extraction. AdaIN [20] was the first to propose the use of channel-wise statistics to modulate feature styles, enabling flexible transfer between arbitrary styles. MixStyle [38] extended this idea to domain generalization in image classification by mixing style statistics across samples to enhance model robustness. For semantic segmentation, SPC-Net [29] models style statistics as prototypes, which guide the model’s generalization to unseen domains based on known style representations. FADA [39] further leverages wavelet transforms to project features into the frequency domain, enabling the decoupling of style and content for more precise modulation. DR-Adapter [40] builds on this idea to align diverse target domain styles with the source domain distribution, thereby boosting performance under few-shot conditions.
Recently, style modulation in object detection remains relatively underexplored, often relying on coarse-grained style bases and simplistic modulation processes. Lopez Rodriguez et al. [41] was the first to apply instance normalization for style unification between source and target domain images, improving feature stability and robustness. TDD [42] utilizes Fourier-based style mapping to translate source-domain images into target-domain styles and employs supervised learning to enhance the detector’s ability to capture target-specific characteristics. However, these methods typically require access to target-domain images, making them less applicable to domain generalization scenarios. To address this issue, FSDA-DETR [9] designed a style projection module to rectify target-domain styles toward the source domain, thereby reducing style shift in few-shot cross-domain detection. LOSA [10] further improved cross-domain adaptability by performing online style projection at both the image and instance levels. In contrast to the above methods that rely on simple storage and updating of style prototypes, our proposed ODS-Adapter constructs a dynamic style memory bank that self-organizes into a set of discriminative style bases, effectively capturing the diversity of training styles, thereby enabling more precise style modulation.
II-D Test-Time Adaptation for Object Detection
Test-Time Adaptation (TTA) is a learning paradigm that aims to update a pre-trained model during inference using unlabeled target-domain data. Tent [43] optimizes batch normalization parameters online by minimizing the entropy loss, thereby enabling the model to adapt to the target domain. DUA [44] eliminates the need for backpropagation and entropy loss, instead achieving adaptation by updating BN statistics in real time. TT-WA [45] dynamically adjusts the affine parameters of the Batch Normalization layers during inference, enabling efficient adaptation to unseen weather conditions. Further, ActMAD [46] models the distribution of entire feature channels to perform more precise feature rectification.
An alternative line of work adopts adaptation through online generated pseudo-labels. MemCLR [47] employs a mean teacher framework that generates pseudo labels to continuously train the detector on the target domain during inference. AMROD [48] introduces a dynamic thresholding mechanism to improve pseudo-label quality. MLFA [49] enhances TTA performance by integrating multi-level feature alignment modules into the paradigm. Differently, we introduce a TTA paradigm designed to address the under-adaptation problem caused by significant discrepancies between training and unseen environments. By continuously absorbing information from the target domain and dynamically updating a memory bank of style representations, our method enables effective adaptation to diverse target domain styles during inference. Since our approach does not rely on generating or using pseudo-labels, it maintains a streamlined, computationally efficient architecture and achieves high adaptation speed.
@@ III Method
Figure 1: Details of (a) the overview of the proposed SA-DETR framework, which adopts DINO as the base detector and incorporates an Online Domain Style Adapter (ODS-Adapter) and an Object-aware Contrastive Learning (OCL) module, and (b) the workflow of the ODS-Adapter. Meanwhile, the memory module follows a test‑time adaptation architecture: it operates during both training and inference to store the statistics of input features and construct multiple style bases, as indicated by the blue arrows. (c) illustrating the projection process during inference. The adapter rectifies the channel wise statistics of the feature maps extracted by the backbone using the stored style bases, thereby mapping the style of the unseen target domain onto that of the training domain.
@@ III-A1 DETR revisit
In our approach, we utilize DINO[50] as the base detector, an advanced DETR-based end-to-end object detector. Specifically, DINO incorporates a CNN backbone for feature extraction and employs an encoder-decoder transformer structure along with a detection head for the final detection predictions. When processing an image, the backbone initially extracts multi-scale feature maps similar to those used in CNN-based detectors. These feature maps are then reshaped into feature tokens that feed into the encoder-decoder module. The encoder further refines the input features by using self-attention. The decoder uses inserted object queries to probe local regions and aggregate instance features from the refined features through cross-attention, generating output embeddings. Finally, these output embeddings are used to compute prediction results through the detection head.
III-A2 Data augmentation
@@
The objective of single-source domain generalization is to train the network with data from a single domain and to generalize its performance to a variety of unseen scenarios. In this task, data augmentation methods are essential to break the limitations of sample diversity. Inspired by previous work[32], we employ spurious correlation generator to augment the source dataset at the image level, ultimately yielding an augmented dataset of equivalent size. Therefore, the input to our method includes both the original and augmented data. Building upon this foundation, we propose the ODS-adapter to leverage known styles for representing unseen ones, thereby overcoming the limitations of augmentation methods in improving performance.
III-A3 Framework
@@
Fig. 1, Our SA-DETR adopts DINO as the base detector and integrates our designed methods, including an ODS-Adapter and an OCL module. Specifically, during training, the network inputs consist of pairs of images, each pair containing an equal number of images from both the source and augmented domains. For a batch of image pairs, the detector first uses a ResNet backbone[51] to extract multi-scale feature maps, and feeds the maps from each layer into our proposed adapter. The adapter leverages a dynamic style memory bank to cache the channel-wise statistics of backbone features and to construct multiple style bases. Then, we insert an OCL module into the encoder during training, which utilizes contrastive learning to encourage the extraction of instance-level domain-invariant features, thereby enhancing the detector’s generalization capabilities. Ultimately, the refined features are input into the subsequent decoder and detection head to obtain detection results, in line with DINO. During inference, the adapter rectifies the channel‑wise statistics of the feature maps extracted by the backbone using the stored style bases, thereby projecting the style of the unseen target domain onto that of the training domain. Notably, to mitigate under‑adaptation arising from large discrepancies between the training and unseen environments, our adapter follows a TTA framework that continuously absorbs target domain information and dynamically updates a memory bank of style bases. Our OCL module is used only during training as a plug in to facilitate the detector’s representation learning. More details on the ODS-Adapter and the OCL module will be presented in Subsections III-B and III-C, respectively.
III-B Online Domain Style Adapter
We present ODS-Adapter that dynamically projects target domain style representations onto the training domain manifold, enabling robust adaptation to diverse unseen scenarios. The adapter first constructs a dynamic style memory bank that self-organizes into multiple discriminative style bases. At inference, it applies a weighted AdaIN projection to achieve fine grained style modulation, thereby enhancing detector robustness on unseen domains. Moreover, ODS-Adapter supports test time adaptation by continually updating the memory bank, effectively mitigating mismatches between training and deployment environments.